The Measure of angle A is twelve more than five times the Measure of Angle B, The angles are Supplementary. Find the Measures of the Angles

Answer by natts9678(5)

Let Angle B = x,
then Angle A = 12 +5x
Given the angles are supplementary, therefore x + 12 + 5x = 180 degrees
==> 6x + 12 = 180
==> 6x = 180 - 12
==> 6x = 168
==> x = 28 degrees
So angle A = 28 degrees and angle B = 152 degrees